# Hyperuniformity and optimal transport of point processes

### v3

@@ 2.2.1 Determinantal processes
where the sum runs over $k$-tuples of pairwise distinct points. Note that not all functions $K$ give rise to a DPP and in particular we will require that $K$ is Hermitian and positive definite; see [8, Section 4.5] for unicity and existence questions. A degenerate example is the unit intensity homogeneous Poisson process for which $\rho_{k}\equiv 1,$ i.e. $K(x,y)=\mathbf{1}_{\{x=y\}}.$ A non-trivial and prominent example is the (unit intensity) infinite Ginibre ensemble on $\mathbb{R}^{2}\approx\mathbb{C}$, defined by
@@
and $\rho_{2}$ only depends on the difference $x-y$, as it turns out the Ginibre ensemble is indeed stationary. It is at the crossroads of combinatorics, statistical physics, and random matrices, as it is also the planar Coulomb gas with inverse temperature $2$, and it also arises as the infinite limit of the process formed by the eigenvalues of a $n\times n$ matrix with i.i.d. standard complex Gaussian entries (see [8]).
@@
A planar stationary determinantal process with kernel $K$ is a $L^{2}$-perturbed lattice (i.e., $\ \mathbf{E}(X^{2})<\infty$ for typical matching distance $X$) if it is hyperuniform, i.e.
@@ 2.2.2 Coulomb gases and other hyperuniform processes
An exception is the zero set $\mu$ of the planar Gaussian entire function, for which the exact fast decay of the RPCM can be inferred from the explicit formulas of the second order correlation measure of [8, Section 3]. Thus combined with Theorem 2, this would yield that such zero sets are $L^{2}$-perturbed lattices; it has actually already been shown in [47] that this is the case, and the authors gave a strong concentration bound on the typical matching distance with an explicit and elegant transport plan to the Lebesgue measure built from the gravitational flow generated by $\mu$. The stable matching of lattice (or some ’nice’ HU point processes) to a Poisson process (or more generally DPPs) of larger intensity produces a hyperuniform point process whose reduced pair correlation satisfies the log integrability condition; see [39, Proposition 8.2] and [44]. But in this case, the construction itself yields a good matching with the lattice.
A particular focus has been put in both the mathematical and physics litterature on 2D Coulomb gases, also called one component plasma (OCP) or Jellium, of which the Ginibre ensemble is a particular case at temperature $1/2$; see [57, 42, 51]. Hyperuniformity of the 2D OCP has been proved by Leblé [42] after having been conjectured for a long time. More precisely he proved that $\sigma(r)\leqslant c\ln(r)^{-0.6}$, but physicists estimate that the second order asymptotic independance is actually much stronger, meaning that the discrepancy is minimal, i.e. $\sigma(r)\sim r^{-1}$, falling in the class I of hyperuniform processes, but it remains for now out of reach for a mathematical treatment.
@@ 2.3 Extension to Random measures
Suppose $\mu$ is a unit intensity stationary random measure. Then $\mathbf{E}\mu(A)=|A|$. Further, relying again on the Campbell-Mecke formula, we can define the RPCM $\beta$ and the diagonal intensity $\lambda_{D}\geq 0$ as follows. Let $g:(\mathbb{R}^{d})^{2}\to\mathbb{R}_{+}$ be a measurable function.
@@
In case $\mu$ is a simple point process, $\lambda_{D}=1$ and often extension of point process results to random measures need to suitably adapt to the fact that $\lambda_{D}\neq 1$. In case the random measure is diffuse (i.e., has not atoms a.s.), then $\lambda_{D}=0$.
@@ 3.1 Fourier-analytic bounds for distance between probability measures
We now state Fourier-analytic bounds for probability measures. These are essentially contained in [13, Proposition 2] and [12, Theorem 1.1]. Similar Fourier-analytic bounds are used in [16] and [15]. See [55] for more applications of such bounds.
@@
$f_{P}(m):=\int_{\Lambda_{1}}e^{i\,m\cdot x}P(\mathrm{d}x),\,\,m\in\mathbb{Z}^{d}.$
For notational convenience, we shall not always indicate that $f_{P}$ is defined over $\mathbb{Z}^{d}$ only. If $P$ is a probability measure and $X$ is a random variable with distribution $P$, we use $f_{X}$ to denote is Fourier-Stieltjes transform $f_{P}$. Also by linearity, we write $f_{P}-f_{Q}$ as $f_{P-Q}$ in case of two probability measures $P,Q$. Also recall from Section 2.1, the notion of transport costs $\widetilde{\mathsf{W}}_{p},\widetilde{\mathsf{C}}_{w}$ under the toroidal distance. Recall that $q:[1,\infty)\to(0,\infty)$ is a non-decreasing function with
@@
Let $Q$ be the uniform probability distribution on $\Lambda_{1}$. There is a constant $c\in(0,\infty)$ (possibly depending on $d$) such that for a probability measure $P$ on $\Lambda_{1}$ and ${t_{0}}>1$,
@@
Suppose $w(x)$ is a continuity modulus on the torus. Then for a constant $c\in(0,\infty)$ (possibly depending on $d$)
@@
$\widetilde{\mathsf{C}}^{2}_{w}(P,Q;\Lambda_{1})$ $\leqslant\,c\,q(t_{0})\,{\sum_{0<\|m\|\leqslant t_{0}}w(\|m\|^{-1})^{2}|f_{P-Q}(m)|^{2}}+cw(t_{0}^{-1})^{2}.$ (21)
In case of $C_{w}$, (20) was proven in [12, Proposition 7.1] (see also [14]) for probability measures supported in $[0,\pi]^{d}$ and (19) in [13, Proposition 2]. Even if the necessary estimates to derive the above bounds are present in these papers, the bounds aren’t stated in the form as above and so we sketch the details now. We will make more exact references to the results we borrow from [12] and [13]. We shall first prove (20), (21), (19) and then (18).
@@
Let $\tilde{P},\tilde{Q}$ the smoothed measures obtained by convoluting $P,Q$ with $\widetilde{t_{0}^{-1}H}$, with the convolution considered in the Torus $\Lambda_{1}$.
We skip $\Lambda_{1}$ from $\tilde{C}_{w}$, $\tilde{W}^{2}_{1}$, $\tilde{W}^{2}_{2}$ for convenience. Estimating that (see also [12, (7.3)])
@@
[…]
Substituting this in (23) and (22) completes the proof of (20). When squaring both sides, we have used that $(a+b)^{2}\leqslant 2a^{2}+2b^{2}$ and absorbed the $2$ into the constant $c$.
To prove (21) (the analogue of [14, (7)]), we observe that in (20), it suffices to choose $q$ defined only on $[1,t_{0}]$. By non-decreasing property, $q(\|m\|)\leqslant q(t_{0})$ and then one can check that $A_{q}^{2}q(t_{0})$ is minimized over $q$ when $q(2^{k})$ is constant for $2^{k}\leqslant t_{0}$. This yields that $A_{q}^{2}q(t_{0})\leqslant c\ln(t_{0})$ and thereby giving us (21).
As for (19), [13, Lemma 1] gives that
@@
using the fact that $H$ has a finite second moment. Now we extend to arbitrary probability measures $P$ by an approximation argument. Observe that by separability of probability measures on a compact Polish space under the $W_{2}$ distance [2, Corollary 8.11], we have that a general probability measure $P$ can be approximated in $\tilde{W}_{2}$ by a sequence of atomic measures $P_{l},l\geq 1$ i.e., $\tilde{W}_{2}(P,P_{l})\to 0$ as $l\to\infty$. Further, weak convergence of $\tilde{P}_{l}$ to $\tilde{P}$ along with quadratic moment convergence (as $\Lambda_{1}$ is compact) gives that $\tilde{W}_{2}(\tilde{P},\tilde{P}_{l})\to 0$ as $l\to\infty$ (see for example, [2, Theorem 8.8]). Now the Lemma follows for arbitrary probability measure $P$ from these two approximations and triangle inequality. ∎
@@
$\tilde{W}^{2}_{2}(\tilde{P},\tilde{Q})$ $\leqslant c\sum_{m\neq 0}\frac{1}{\|m\|^{2}}|f_{\tilde{P}-\tilde{Q}}(m)|^{2}\leqslant c^{\prime}\sum_{m\neq 0}\frac{1}{\|m\|^{2}}|f_{P}(m)|^{2}$
@@ 3.2 Proof of Theorems 3 and 4
Recall that $\mu$ is a stationary point process with unit intensity. Set ^$\hat{\mathcal{L}}_{n}:=\frac{1}{{n}(2\pi)^{d}}\mathcal{L}1_{\Lambda_{n}}$. The proof is based on Theorem 6 with the rescaled sample on $\Lambda_{1}$
$\hat{\mu}_{n}=N^{-1}\mu(1_{\Lambda_{n}}n^{1/d}\cdot)$
where $N=\mu(\Lambda_{n})$ and set $\hat{\mu}_{n}=\delta_{0}$ if $N=0$. Observe that $\hat{\mu}_{n},\hat{\mathcal{L}_{n}}$ are probability measures. Recall also the (not-rescaled) renormalised sample
$\tilde{\mu}_{n}=\frac{n}{N}\mu 1_{\Lambda_{n}}.$
with mass $n$ and again $\tilde{\mu}_{n}=n\delta_{0}$ if $N=0$. Traditionally, the scattering intensity of $\mu_{n}$ is defined as
$S_{n}(k)=\frac{1}{N}|\mu(1_{\Lambda_{n}}e^{ik\cdot})|^{2}$ (26)
and $S_{n}\equiv 0$ if $\mu=0.$ Thus for $m\in\mathbb{Z}^{d}$
$|f_{\hat{\mu}_{n}}(m)|^{2}=\frac{1}{N^{2}}|\mu(1_{\Lambda_{n}}e^{imn^{-1/d}\cdot})|^{2}=\frac{1}{N}S_{n}(mn^{-1/d}),$ (27)
and set $f_{\hat{\mu}_{n}}(m)=1$ if $N=0$.
Recall the definitions of $S,\mathsf{b}_{n},\epsilon(t)$ from (6),(8) and (9) respectively. To prove Theorems 3 and 4, we need some estimates on variance and scattering intensity. The first lemma is a quantitative refinement of [45, Proposition 2] and the second borrows a trick from [27], probably well known in the physics litterature. It quantifies the rate of convergence of the scattering intensity to the structure factor; see [20, Section 5], [38, Section 3].
@@
If $\beta$ is integrable,
$\text{Var}(N)=n\Big{(}\beta(\mathbb{R}^{d})+1+O(\varepsilon(n^{-1/d})\Big{)}.$ (28)
Furthermore, if $\beta(A)\leqslant 0$ for $A\subset\mathbb{R}^{d}$,
$\text{Var}(N)\geqslant n\Big{(}1+\beta(\mathbb{R}^{d})+\varepsilon(n^{-1/d})\Big{)}.$
@@
$\mathbf{E}\left(\frac{n}{N}S_{n}(k)\right)=S(k)+\frac{\text{Var}(N)}{n}+O(\varepsilon(n^{-1/d})).$ (29)
@@
Let us conclude the proofs of Theorems 3 and 4 before proving the lemmata. The proof ideas are to use Theorem 6 by rescaling $\tilde{\mu}_{n}$ and ~$\tilde{\mathcal{L}}_{n}$ to probability measures $\hat{\mu}_{n},\hat{\mathcal{L}}_{1}$ on $\Lambda_{1}$ and comparing the same to the costs on $\Lambda_{n}$.
@@
We again skip $\Lambda_{n}$ from $\tilde{W}_{p},p\in[0,1]\cup\{2\}$ for convenience. Fix $t_{0}=n^{1/d}$. By (27) and Lemma 9, $\mathbf{E}(|f_{\hat{\mu}_{n}}(m)|^{2})\leqslant c\mathsf{b}_{n}n^{-1}$ for $\|m\|\leq t_{0}$. So using Theorem 6 (in fact (19) and (18)) yields for $p=1,2$
@@
$\leqslant$ $\,cn^{-1}\mathsf{b}_{n}{\sum_{k=1}^{n^{1/d}}\frac{k^{d-1}}{k^{2}}}+cn^{-2/d}$
[…]
Fix $p\in(0,1)$. We can derive three different bounds in this case and then choose the best possible. Firstly, two successive uses of Hölder’s inequality and noting that $\hat{\mu_{n}}$ and ^$\hat{\mathcal{L}}_{1}$ are probability measures, gives the following bound
@@
$\leqslant$ $\,cn^{-1}(\mathsf{b}_{n}+1){\sum_{k=1}^{n^{1/d}}\,q(k)\,k^{d-1-2p}}+n^{-2p/d}.$ (32)
Finally, using (21), we derive
$\mathbf{E}\widetilde{\mathsf{W}}_{p}^{2p}(\hat{\mu}_{n},\hat{\mathcal{L}}_{1})\leqslant\,cn^{-1}(\mathsf{b}_{n}+1)\,\ln(n)\,{\sum_{k=1}^{n^{1/d}}\,k^{d-1-2p}}+n^{-2p/d}.$ (33)
Now using (32) for $d=1,p>1/2$ with $q(x)=\ln(2x)\ln\ln(3x)^{\gamma}$ for $\gamma>1$ (which is summable in this case), using (33) for $d=1,p=1/2$ and (31) in the remaining cases, we obtain that
@@
Going back to large scale: For $p>0,$ let $\hat{M}_{n}$ be a $\widetilde{{\mathsf{W}}}_{p}-$optimal transport plan between $\hat{\mu}_{n}$ and ^$\hat{\mathcal{L}}_{1}$, and $M_{n}(\mathrm{d}x,\mathrm{d}y):=n\hat{M}_{n}(n^{-1/d}\mathrm{d}x,n^{-1/d}\mathrm{d}y)$ is a ${\mathsf{W}}_{p}$ transport plan between $\tilde{\mu}_{n}$ and ~$\tilde{\mathcal{L}}_{n}$. Note that $M_{n}$ may not necessarily be optimal. Then we can derive
@@
Also, by the HU property, and the fact that
$|e^{-ik\cdot x}-1|\leqslant c\min\{|k\cdot x|;1\}$
$S(k)=$ $1+\int e^{-ik\cdot x}\beta(\mathrm{d}x)=\int(e^{-ik\cdot x}-1)\beta(\mathrm{d}x)$
$|S(k)|\leqslant$ $\int\min(1,|k\cdot x|)|\beta|(\mathrm{d}x)\leqslant\varepsilon(\|k\|),$
@@
${\mathbf{E}\widetilde{\mathsf{W}}_{2}^{2}(\tilde{\mu}_{n},\tilde{\mathcal{L}}_{n})}\leqslant c\,n\left({1+\int_{n^{-1/d}}^{1}\varepsilon(r)r^{d-3}\mathrm{d}r}\right),$
@@
$|\int_{0}^{1}\epsilon(r)r^{d-3}\mathrm{d}r|=\int\mathbf{1}[|x|>1]\int_{0}^{1}\min(1,r|x|)r^{d-3}\mathrm{d}r|\beta|(\mathrm{d}x)<\infty.$
So, we have that $\alpha_{2}(n)\leqslant cn$ in (15) where we have chosen $c_{0}=1$. Now the claim follows from item (1) of the theorem.
(3).   Consider (15) with $c_{0}=1$. We shall bound $\int_{n^{-1}}^{1}\epsilon(r)r^{-2}\mathrm{d}r$ as follows. First observe that
[…]
So integrating with $|\beta|(\mathrm{d}x)$, we derive that
[…]
$\leq c\ln(n),$
where in the last inequality we have used $\int|x||\beta|(\mathrm{d}x)<\infty$ to deduce that $n|\beta|(|x|\geq n)\to 0$ and the finiteness of the last integral. Now again by Item (1) of the theorem, the claim follows. ∎
@@
$0<c\leqslant\frac{\gamma_{n}(x)}{n\min\{1;\|x\|n^{-1/d}\}}\leqslant C<\infty$ (36)

### v4

@@ 2.2.1 Determinantal processes
where the sum runs over $k$-tuples of pairwise distinct points. Note that not all functions $K$ give rise to a DPP and in particular we will require that $K$ is Hermitian and positive definite; see [9, Section 4.5] for unicity and existence questions. A degenerate example is the unit intensity homogeneous Poisson process for which $\rho_{k}\equiv 1,$ i.e. $K(x,y)=\mathbf{1}_{\{x=y\}}.$ A non-trivial and prominent example is the (unit intensity) infinite Ginibre ensemble on $\mathbb{R}^{2}\approx\mathbb{C}$, defined by
@@
and $\rho_{2}$ only depends on the difference xyx-y, as it turns out the Ginibre ensemble is indeed stationary; see for example [6, Proposition 5.6.6]. It is at the crossroads of combinatorics, statistical physics, and random matrices, as it is also the planar Coulomb gas with inverse temperature $2$, and it also arises as the infinite limit of the process formed by the eigenvalues of a $n\times n$ matrix with i.i.d. standard complex Gaussian entries (see [9]).
@@
[…]
@@ 2.2.2 Coulomb gases and other hyperuniform processes
An exception is the zero set $\mu$ of the planar Gaussian entire function, for which the exact fast decay of the RPCM can be inferred from the explicit formulas of the second order correlation measure of [9, Section 3]. Thus combined with Theorem 2, this would yield that such zero sets are $L^{2}$-perturbed lattices; it has actually already been shown in [49] that this is the case, and the authors gave a strong concentration bound on the typical matching distance with an explicit and elegant transport plan to the Lebesgue measure built from the gravitational flow generated by $\mu$. The stable matching of lattice (or some ’nice’ HU point processes) to a Poisson process (or more generally DPPs) of larger intensity produces a hyperuniform point process whose reduced pair correlation satisfies the log integrability condition; see [41, Proposition 8.2] and [47]. But in this case, the construction itself yields a good matching with the lattice.
A particular focus has been put in both the mathematical and physics literature on 2D Coulomb gases, also called one component plasma (OCP) or Jellium, of which the Ginibre ensemble is a particular case at temperature $1/2$; see [61, 45, 54]. Hyperuniformity of the 2D OCP has been proved by Leblé [45] after having been conjectured for a long time. More precisely he proved that $\sigma(r)\leqslant c\ln(r)^{-0.6}$, but physicists estimate that the second order asymptotic independance is actually much stronger, meaning that the discrepancy is minimal, i.e. $\sigma(r)\sim r^{-1}$, falling in the class I of hyperuniform processes, but it remains for now out of reach for a mathematical treatment.
@@ 2.3 Extension to Random measures
Suppose $\mu$ is a unit intensity stationary random measure. Then $\mathbf{E}\mu(A)=\mathcal{L}(A)$. Further, relying again on the Campbell-Mecke formula, we can define the RPCM $\beta$ and the diagonal intensity $\lambda_{D}\geq 0$ as follows. Let $g:(\mathbb{R}^{d})^{2}\to\mathbb{R}_{+}$ be a measurable function.
@@
In case $\mu$ is a simple point process, $\lambda_{D}=1$ and often extension of point process results to random measures need to suitably adapt to the fact that $\lambda_{D}\neq 1$. In case the random measure is diffuse (i.e., has not atoms a.s.), then $\lambda_{D}=0$; for example, see [43].
@@ 3.1 Fourier-analytic bounds for distance between probability measures
We now state Fourier-analytic bounds for probability measures. These are essentially contained in [14, Proposition 2] and [13, Theorem 1.1]. Similar Fourier-analytic bounds are used in [17] and [16]. See [58] for more applications of such bounds.
@@
$f_{P}(m):=\int_{\Lambda_{1}}e^{-i\,m\cdot x}P(\mathrm{d}x),\,\,m\in\mathbb{Z}^{d}.$
For notational convenience, we shall not always indicate that $f_{P}$ is defined over $\mathbb{Z}^{d}$ only. $P$ is a probability measure and $X$ is a random variable with distribution $P$, we use $f_{X}$ to denote is Fourier-Stieltjes transform $f_{P}$. Also by linearity, we write $f_{P}-f_{Q}$ as $f_{P-Q}$ in case of two probability measures $P,Q$. Also recall from Section 2.1, the notion of transport costs $\widetilde{\mathsf{W}}_{p},\widetilde{\mathsf{C}}_{w}$ under the toroidal distance. Recall that $q:[1,\infty)\to(0,\infty)$ is a non-decreasing function with
@@
Let $Q$ be the uniform probability distribution on $\Lambda_{1}$. There is a constant $c\in(0,\infty)$ (possibly depending on $d$) such that for any probability measure $P$ on $\Lambda_{1}$ and all ${t_{0}}>1$, we have that
@@
Suppose $Q$ is any probability distribution on $\Lambda_{1}$ and $w$ is a continuity modulus on the torus. Then for a constant $c\in(0,\infty)$ (possibly depending on $d$) such that for any probability measure $P$ on $\Lambda_{1}$ and all ${t_{0}}>1$, we have that
@@
$\widetilde{\mathsf{C}}^{2}_{w}(P,Q;\Lambda_{1})$ $\leqslant\,c\,\ln(t_{0})\,{\sum_{0<\|m\|\leqslant t_{0}}w(\|m\|^{-1})^{2}\,|f_{P-Q}(m)|^{2}}+c\,w(t_{0}^{-1})^{2}.$ (24)
In case of $C_{w}$, (23) was proven in [13, Proposition 7.1] (see also [15]) for probability measures supported in $0,\pi]^{d}$ and (22) in [14, Proposition 2]. Even if the necessary estimates to derive the above bounds are present in these papers, the bounds aren’t stated in the form as above and so we sketch the details now. We will make more exact references to the results we borrow from [13] and [14]. Note that in (21), $f_{P}=f_{P-Q}$ as $f_{Q}(m)=0,m\in\mathbb{Z}^{d}$ for uniform probability measure on $\Lambda_{1}$.
We shall prove (23), (24), (22) and (21), in that order.
@@
Let $\tilde{P},\tilde{Q}$ the smoothed measures obtained by convoluting $P,Q$ with $\widetilde{t_{0}^{-1}H}$ (the projection of $t_{0}^{-1}H$ onto $\Lambda_{1}$), with the convolution considered in the Torus $\Lambda_{1}$.
We skip $\Lambda_{1}$ from $\tilde{C}_{w}$, $\tilde{W}^{2}_{1}$, $\tilde{W}^{2}_{2}$ for convenience. Estimating via the dual representation that (see also [13, (7.3)])
@@
$|f_{\tilde{P}}(m)-f_{\tilde{Q}}(m)|$ $\leqslant|f_{P-Q}(m)|\times|f_{t_{0}^{-1}H}(m)|\leqslant|f_{P-Q}(m)|\,{\mathbf{1}[\|m\|\leqslant t_{0}]},$ (27)
where we have also used that the support of $f_{H}$ is in the unit ball. Substituting this in (26) and (25) completes the proof of (23). When squaring both sides, we have used that $(a+b)^{2}\leqslant 2a^{2}+2b^{2}$ and absorbed the $2$ into the constant $c$.
To prove (24) (the analogue of [15, (7)]), we observe that in (23), it suffices to choose $q$ defined only on $1,t_{0}]$. By non-decreasing property, $q(\|m\|)\leqslant q(t_{0})$ and then one can check that $A_{q}^{2}q(t_{0})$ is minimized over $q$ when $q(2^{k})$ is constant for $2^{k}\leqslant t_{0}$. This yields that $A_{q}^{2}q(t_{0})\leqslant c\ln(t_{0})$ and thereby giving us (24).
As for (22), [14, Lemma 1] gives that
@@
using the fact that $H$ has a finite second moment. Now we extend to arbitrary probability measures $P$ by an approximation argument. Observe that by separability of probability measures on a compact Polish space under the $W_{2}$ distance [3, Corollary 8.11], we have that a general probability measure $P$ can be approximated in $\tilde{W}_{2}$ by a sequence of atomic measures $P_{l},l\geq 1$ i.e., $\tilde{W}_{2}(P,P_{l})\to 0$ as $l\to\infty$. Further, weak convergence of $\tilde{P}_{l}$ to $\tilde{P}$ along with quadratic moment convergence (as $\Lambda_{1}$ is compact) gives that $\tilde{W}_{2}(\tilde{P},\tilde{P}_{l})\to 0$ as $l\to\infty$ (see for example, [3, Theorem 8.8]). Now the Lemma follows for arbitrary probability measure $P$ from these two approximations and triangle inequality. ∎
@@
[…]
@@ 3.2 Proof of Theorems 3 and 4
Recall that $\mu$ is a stationary point process with unit intensity. Set $\hat{\mathcal{L}}_{1}:=\frac{1}{(2\pi)^{d}}\mathcal{L}\mathbf{1}_{\Lambda_{1}}$, the uniform probability measure on $\Lambda_{1}$. The proof is based on Theorem 6 with the rescaled sample on $\Lambda_{1}$
$\hat{\mu}_{n}=N^{-1}\mu(\mathbf{1}_{\Lambda_{n}}n^{1/d}\cdot)$
where $N=\mu(\Lambda_{n})$ and set $\hat{\mu}_{n}=\delta_{0}$ if $N=0$. Observe that $\hat{\mu}_{n},\hat{\mathcal{L}}_{1}$ are probability measures on $\Lambda_{1}$. Recall also the (not-rescaled) renormalised sample
$\tilde{\mu}_{n}=\frac{n}{N}\mu\mathbf{1}_{\Lambda_{n}}.$
with mass $n$ and again $\tilde{\mu}_{n}=n\delta_{0}$ if $N=0$. The scattering intensity of $\mu_{n}=\mu(\mathbf{1}_{\Lambda_{n}}\cdot)$ is defined as
$S_{n}(k)=\frac{1}{N}\left|\int_{\Lambda_{n}}e^{-i\,k\cdot x}\mu(\mathrm{d}x)\right|^{2}$ (29)
and $S_{n}\equiv 0$ if $N=0.$ Thus for $m\in\mathbb{Z}^{d}$
$|f_{\hat{\mu}_{n}}(m)|^{2}=\frac{1}{N^{2}}\left|\int_{\Lambda_{n}}e^{-i\,n^{-1/d}\,m\cdot x}\mu(\mathrm{d}x)\right|^{2}=\frac{1}{N}S_{n}(mn^{-1/d}),$ (30)
and set $f_{\hat{\mu}_{n}}(m)=1$ if $N=0$. The scattering intensity can be seen as an empirical estimate of the structure factor $S$. More precise convergence rates are given below and are useful for our proofs.
Recall the definitions of $S,\mathsf{b}_{n},\varepsilon(t)$ from (8),(11) and (12) respectively. To prove Theorems 3 and 4, we need some estimates on variance and scattering intensity. The first lemma is a quantitative refinement of [48, Proposition 2] and the second borrows a trick from [30], probably well known in the physics literature. It quantifies the rate of convergence of the scattering intensity to the structure factor; see also [21, Theorem 5.1] for a more general weak convergence result.
@@
Let $\pi_{d}:=(2\pi)^{d}$. If $\beta$ is integrable,
$\text{Var}(N)=n\pi_{d}\Big(\beta(\mathbb{R}^{d})+1+O(\varepsilon(n^{-1/d})\Big).$ (31)
Furthermore, if $\beta(A)\leqslant 0$ for all $A\subset\mathbb{R}^{d}$,
$\text{Var}(N)\geqslant n\pi_{d}\Big(1+\beta(\mathbb{R}^{d})+\varepsilon(n^{-1/d})\Big).$
@@
$\mathbf{E}\left(\frac{\pi_{d}n}{N}S_{n}(k)\right)=S(k)+\frac{\text{Var}(N)}{\pi_{d}n}+O(\varepsilon(n^{-1/d})).$ (32)
@@
Let us conclude the proofs of Theorems 3 and 4 before proving the lemmata. The proof uses Theorem 6 by rescaling $\tilde{\mu}_{n}$ and $\tilde{\mathcal{L}}_{n}$ to probability measures $\hat{\mu}_{n},\hat{\mathcal{L}}_{1}$ on $\Lambda_{1}$ and comparing the same to the costs on $\Lambda_{n}$. Note that for all $m\in\mathbb{Z}^{d}$, $f_{\hat{\mathcal{L}}_{1}}(m)=0$ as it is the uniform probability measure on $\Lambda_{1}$ and so $f_{\hat{\mu}_{n}-\hat{\mathcal{L}}_{1}}(m)=(f_{\hat{\mu}_{n}}-f_{\hat{\mathcal{L}}_{1}})(m)=f_{\hat{\mu}_{n}}(m)$.
@@
We again skip $\Lambda_{n}$ from $\tilde{W}_{p},p\in[0,1]\cup\{2\}$ for convenience. Fix $t_{0}=n^{1/d}$. By (30) and Lemma 9, $\mathbf{E}(|f_{\hat{\mu}_{n}}(m)|^{2})\leqslant c\mathsf{b}_{n}n^{-1}$ for $\|m\|\leq t_{0}$. So using Theorem 6 (in fact (22) and (21) with $f_{\hat{\mathcal{L}}_{1}}(m)=0$) yields for $p=1,2$
@@
$\leqslant$ $\,cn^{-1}\mathsf{b}_{n}{\sum_{j=1}^{n^{1/d}}\frac{j^{d-1}}{j^{2}}}+cn^{-2/d}$
$\leqslant$ $\,c\mathsf{b}_{n}\begin{cases}n^{-1}$ if $d=1\\ n^{-1}{\ln(n)}$ if $d=2\\ n^{-1}{(n^{1/d})^{d-2}}=n^{-2/d}$ if $d\geqslant 3\end{cases}$
Fix $p\in(0,1)$. We can derive three different bounds in this case and then choose the best possible. Firstly, two successive uses of Hölder’s inequality and noting that $\hat{\mu}_{n}$ and $\hat{\mathcal{L}}_{1}$ are probability measures, gives the following bound
@@
$\leqslant$ $\,cn^{-1}(\mathsf{b}_{n}+1){\sum_{j=1}^{n^{1/d}}\,{q(j)}\,j^{d-1-2p}}+n^{-2p/d}.$ (35)
Finally, using (24) and that $q(t_{0})\leq c\ln(n)$, we derive
$\mathbf{E}\widetilde{\mathsf{W}}_{p}^{2p}(\hat{\mu}_{n},\hat{\mathcal{L}}_{1})\leqslant\,cn^{-1}(\mathsf{b}_{n}+1)\,\ln(n)\,{\sum_{j=1}^{n^{1/d}}\,j^{d-1-2p}}+n^{-2p/d}.$ (36)
Now using (35) for $d=1,p>1/2$ with $q(x)=\ln(2x)\ln\ln(3x)^{\gamma}$ for $\gamma>1$ (which is summable in this case), using (36) for $d=1,p\leq 1/2$ and (34) in the remaining cases, we obtain that
@@
Going back to large scale: For $p>0,$ let $\hat{M}_{n}$ be a $\widetilde{{\mathsf{W}}}_{p}$-optimal transport plan between $\hat{\mu}_{n}$ and $\hat{\mathcal{L}}_{1}$, and $M_{n}(\mathrm{d}x,\mathrm{d}y):=n\hat{M}_{n}(n^{-1/d}\mathrm{d}x,n^{-1/d}\mathrm{d}y)$ is a ${\mathsf{W}}_{p}$ transport plan between $\tilde{\mu}_{n}$ and $\tilde{\mathcal{L}}_{n}$. Note that $M_{n}$ may not necessarily be optimal. Then we can derive
@@
Also,the HU property and
$|e^{-ik\cdot x}-1|\leqslant c\min(1,|k\cdot x|)\,\,\textrm{and}\,\,S(k)=$ $1+\int e^{-ik\cdot x}\beta(\mathrm{d}x)=\int\,(\,e^{-ik\cdot x}-1\,)\beta(\mathrm{d}x)$
gives us that
$|S(k)|\leqslant$ $\,c\,\int\min(1,|k\cdot x|)|\beta|(\mathrm{d}x)\leqslant\,c\,\varepsilon(\|k\|),$
@@
${\mathbf{E}\widetilde{\mathsf{W}}_{2}^{2}(\tilde{\mu}_{n},\tilde{\mathcal{L}}_{n})}\leqslant c\,n\left({1+\int_{n^{-1/d}}^{c_{0}}\varepsilon(r)r^{d-3}\mathrm{d}r}\right)=\alpha_{2}(n),$ (39)
@@
$\int_{0}^{1}\varepsilon(r)r^{d-3}\mathrm{d}r=\int|x|\mathbf{1}[|x|<1]|\beta|(\mathrm{d}x)+\int\mathbf{1}[|x|>1]\int_{0}^{1}\min(1,r|x|)r^{d-3}\mathrm{d}r|\beta|(\mathrm{d}x)<\infty.$
So, we have that $\alpha_{2}(n)\leqslant cn$ in (18) where we have chosen $c_{0}=1$. Now the claim follows from item (1) of the theorem i.e., (39). ∎
@@
$0<c\leqslant\frac{\gamma_{n}(x)}{n\min(1,\|x\|n^{-1/d})}\leqslant C<\infty$ (40)